Skip to content

By Darrin Crone on

Sir Nigel Gresley overhaul – Update 25

The latest update from the workshop, where a historic locomotive is being restored.

Week commencing 5 February 2018

More work has been done on cleaning and descaling the coal space in the tender. A considerable amount of scale has been removed and barrowed out to the skip. The GSMR rack overhaul continues with primer now being applied.

The tender coal space and the top of the corridor is being descaled. A large amount of material has been removed.

The dye pen inspection of the coupling rods has continued this week and the outside connecting rods have been similarly treated. No defects have been found. The inside connecting rod was cleaned up and measured. The small end (gudgeon) pins and the gradient pins were cleaned and examined. A refurbishment program for the pins and the inside connecting rod is now being devised.

The inspection of the coupling rods has continued, with the ends inspected for cracks.
The little end (or gudgeon) pins have been thoroughly prepared for examination. They have now been measured for wear.
The inside connecting rod has been retrieved from store, cleaned up and a start made on measuring it as part of its examination.

Work continued on the drain cock linkage, as we worked on the cross shaft located under the trailing edge of the middle cylinder. This required a couple of the pin holes to be attended to. One was remade while the material on the edge of the link was built up.

The inaccessible nut on one of the cylinder flange bolts was driven home this week after the nut was specially machined to accept a punch. The assessment of the hornstay bolting continued with the removal and examination of the horizontal hornstay bolts. As with the vertical bolting, there is plenty of evidence of fitter examinations with damage from hammers. Quotations are being sought for new nuts and bolts.

We continued tidying up the wheelset paint line along the tyre edges. They look very good. On Saturday our 007 Gang of junior volunteers checked the identity of the bogie wheels to ensure they are the right way round to be put back in to the bogie frames. Each journal was measured with a micrometer and it was found necessary to rotate the trailing set. All wheelsets are now in the correct orientation.

The tyre edges on the bogie wheels have been cleaned off, the trailing set has been turned to get it the right way round and the protective covers have been removed from the journals. The wheelsets sit on the wheel drop in the National Railway Museum workshop, originally part of York North depot.

The dimensions and calculations for the coupled wheel axleboxes’ machining dimensions are coming together and have been checked. Final measurements and checks of the coupled wheel journals have been carried out and final dimensions will be available for checking soon.

The LNER blow down cock was given a final examination and the packings were given an additional squeeze. Another boiler component, the regulator to the main steam pipe clamp, was retrieved from store this week and descaled as it is now required at Llangollen. Work on the ashpan continues with more new platework put into the rear corners and other areas prepared. The reaming of the superheater header continued this week and we are now more than half way.

The ashpan back right back corner has received further new plate.
The clamp that pulls the regulator valve onto the internal steam pipe inside the boiler barrel was retrieved from store this week and prepared to be sent to Llangollen.

The bench measurements of the slidebars are now complete and the results have been tabulated and sent to our CME. The slidebars will be subject to further examination for defects and will then be refitted. Work started this week on preparing the brackets on the loco to accept the slidebars.

The lubrication pipework for the oil boxes that feed the slidebars and the areas around them was retrieved from store this week and a start made on assessing their condition has begun. A number of the pipes have now been annealed.

To allow pipe fitting to progress at the cab end we need the cab floor plate in position, and so this was refitted this week. At the moment the floor plate is secured by temporary screws as most of the old screws and bolts were beyond further use. Those that can be reused have been refurbished. The floor plate went in very easily with all the shims previously removed refitted in the correct location using the records made during removal. New bolting is on order.

The floor plate is lowered into position by Roger Turnbull and Ian Seales.

The cab upstand platework that fastens to the floor plate and supports the cab floor has been cleaned and descaled. Some has had to be scrapped while the largest that goes across the width of the cab above the bufferbeam has been primed. New steel has been ordered to replace the pieces that have been condemned.

At Llangollen work progresses on the tube plate and the regulator set up. This photo shows inside the stuffing box on the backhead with the reach rod that goes through the boiler to the regulator valve.

This is the 25th update. You can read the previous instalments here.

Leave a comment

Your email address will not be published. Required fields are marked *